OpenStudy (anonymous):

Indicate which property is illustrated in Step 5. Step 1 9x + 11 + 12x + 10 = (9x + 11) + (12x + 10) Step 2 = 9x + (11 + 12x) + 10 Step 3 = 9x + (12x + 11) + 10 Step 4 = (9x + 12x) + (11 + 10) Step 5 = (9 + 12)x + (11 + 10) Step 6 = 21x + 21 associative commutative arithmetic fact distributive

Directrix (directrix):

The question is which property allows a person to go from: (9x + 12x) + (11 + 10) to (9 + 12)x + (11 + 10)

Directrix (directrix):

Notice that the (11 + 10) portion did not change. So, the task lies in determining how to get from: (9x + 12x) to (9 + 12)x @Starr_Dynasty Any ideas? Does that look familiar?

OpenStudy (anonymous):

i need the property @Directrix i put it at the bottom

Directrix (directrix):

@Starr_Dynasty I know you need the property, the answer. I am trying to ask you about it. I see the answer and am trying to help you see it.

OpenStudy (anonymous):

no it really doesnt look familiar but im appreciating the help @Directrix

Directrix (directrix):

@Starr_Dynasty Okay, click on the attachment and think if it looks like what has happened in moving from step 4 to step 5. (9x + 12x) = (9 + 12)x

Directrix (directrix):

(9x + 12x) = (9 + 12)x ---> The x has been distributed out of the 9x + 12x. We usually don't think of the distributive property in that way.

Directrix (directrix):

Question?